A man stands on the roof of a 15.0-m-tall building and throws a rock with a velocity of magnitude 30 m/s at an angle of 33° above the horizontal. You can ignore air resistance. Calculate. A) the maximum height above the roof reached by the rock; B) the magnitude of the velocity of the rock just before it strikes the ground; and C) the horizontal range from the base of the building to the point where the rock strikes the ground.


1
Expert's answer
2021-04-26T17:11:45-0400


(a) At maximum height, i.e. at point P:

Vertical component of velocity "v_y =0"

Initial horizontal speed

"x_{0x}=v_0cosθ \\

= 30 \times cos 33º \\

= 25.2 \;m/s"

Initial vertical speed

"v_{0y}=v_0sinθ \\

= 30 \times sin 33º \\

= 16.3 \;m/s"

Let H be the maximum height reached by the rock, as measured from the roof.

"v^2_y=v^2_{0y}-2gH"

Maximum height

"H= \frac{v^2_{0y}}{2g} \\

= \frac{16.3}{2 \times 9.8} \\

= 13.6 \;m"

(b) Magnitude of velocity of the rock just before it strikes the ground

"v = \sqrt{v^2_x+v^2_y} \\

v_x = v_{0x} \\

= 25.2 \;m/s"

Time of flight (t) is given by

"-h = v_{0y}t- \frac{1}{2}gt^2 \\

4.9t^2 -16.3t -15 =0 \\

t = \frac{16.3± \sqrt{16.3^2 + 4 \times 4.9 \times 15}}{9.8} \\

= \frac{16.3±23.7}{9.8}"

= 4.08 s (ignoring the negative value)

"v_y = 16.3 – 9.8 \times 4.08 \\

= -23.7 \;m/s \\

v = \sqrt{v^2_x + v^2_y} \\

= \sqrt{25.2^2 + (-23.7)^2} \\

= 34.6 \;m/s"

(c) Required horizontal distance "(x) = v_{0x}t"

"= 25.2 \times 4.08 \\

= 102.8 \;m"